# From Laboratory to Real-World Applications: Benchmarking Agentic Code Reasoning at the Repository Level

### v1

@@ 2.2 Code Reasoning
Recent benchmarks like CRUXEval (Gu et al., 2024) and REval (Chen et al., 2024) pioneered execution reasoning but are largely confined to isolated, “lab-scale” code snippets. While tools like CORE (Xie et al., 2025) incorporate static analysis tasks to evaluate reasoning, they lack the structural depth and code breadth (e.g., deep inheritance, cross-module imports) characteristic of large-scale projects. RepoReason bridges this gap by applying deep execution simulation to massive repository scopes, challenging models to maintain logical consistency across extensive file systems.
@@ 2.3 Agentic Software Engineering
At the highest level, SWE-bench (Jimenez et al., 2024) evaluates agents on resolving GitHub issues. However, as an end-to-end evaluation, it acts as a “black-box” indicator of success or failure. RepoReason complements this by targeting the cognitive root causes of failure. By decoupling reasoning from code editing and utilizing fine-grained metrics (ESV, MCL, DFI), it provides the “white-box” diagnostic insights required to pinpoint an agent’s specific bottlenecks in context handling or multi-source integration.
@@ 3 RepoReason: Design and Construction
In this section, we delineate the methodological framework of RepoReason, conceptualizing the construction pipeline as a rigorous five-phase defense system. Our objective is to transform authentic software logic into “white-box” diagnostic tasks while systematically neutralizing systemic vulnerabilities: spurious correlation, rote memorization, and logical superficiality. As illustrated in Figure 1, the pipeline orchestrates sequential gates: (1) Repository Curation to ensure environmental fidelity; (2) Structural Filtering via runtime traces to guarantee logical depth; (3) Execution-Driven Mutation to eliminate parametric leakage; (4) Task Instantiation governed by a deterministic value protocol for stability; and (5) Diagnostic Evaluation establishing fine-grained cognitive profiles. This integration ensures that each task instance necessitates authentic repository-level reasoning and robust state tracking, effectively precluding resolution via shallow pattern matching.
@@ 3.1 Design Philosophy: Assertions as Semantic Anchors
Spurious Correlation in Raw Traces: A straightforward approach might involve randomly querying variable values at specific locations (e.g., “What is the value of variable $x$ at line 10?”). However, this “Raw Trace Querying” is fundamentally flawed for evaluating deep reasoning. It suffers from shortcut learning, as models can often retrieve answers through shallow static analysis of nearby assignment statements, bypassing the need for genuine execution simulation.
@@ 3.2 Foundation: Real-world Repository Curation
To ensure that our “Semantic Anchors” are grounded in reality, we curated a set of mature, actively maintained Python repositories. This stage ensures the fidelity of our benchmark, providing the complex, cross-file dependency structures that synthetic datasets lack.
@@ 3.6 Cognitive Diagnostic Framework
To quantify the intrinsic reasoning complexity of a task, we perform dynamic program slicing (Agrawal and Horgan, 1990) rooted at the assertion to obtain the “Minimal Causal Computational Subgraph” along the actual execution path. Dynamic cross-function backward slicing is obtained by tracing back from the assertion along data and control dependencies within a recorded execution trace to isolate the minimal set of statements that causally influence the observed result. Based on this subgraph, we characterize reasoning intensity using three mutually orthogonal metrics rooted in cognitive load theory (Sweller, 1988).
@@ 4.2 Overall Performance: Accuracy across Repositories and Difficulty Levels
1. Explicit Algorithmic Patterns (Cachetools and Toolz): In repositories such as cachetools and toolz, models demonstrate robust and consistent performance, with mean accuracies ranging from 78.00% to 82.00%. These libraries primarily rely on standard algorithmic paradigms—such as cache eviction and functional data flows—characterized by high logical locality and linear execution transitions. The results indicate that LLMs have effectively mastered these procedural patterns, which currently define the “Cognitive Comfort Zone” for agentic reasoning.
2. Implicit Metaprogramming Logic (Attrs and Jinja2): attrs and jinja2 exhibit a significant performance stratification. Taking attrs as an example, which relies heavily on decorators for implicit class generation (e.g., the dynamic injection of __init__ methods), the performance gap between Claude (72.41%) and Kimi-K2 (43.10%) reveals a fundamental divergence in the internalization of the Python dynamic object model. Frontier models demonstrate the capacity to model runtime behaviors that remain “invisible” as explicit code lines in static text, whereas other models show a marked deficiency when logic is generated programmatically rather than defined literally.
3. High-Entropy Symbolic and Structural Reasoning (SymPy and NetworkX): sympy and networkx represent the primary “Reasoning Ceiling” for current agents. Symbolic transformations in sympy involve abstract, non-linear logical transitions that frequently cause reasoning chains to deviate. Concurrently, networkx imposes a high demand on Structured Working Memory to maintain complex graph topologies. Claude’s superior performance in networkx (79.77%) confirms its advanced capability to maintain accurate “state snapshots” during complex traversals without experiencing disorientation.

### v2

@@ 2.2 Code Reasoning
Recent benchmarks like CRUXEval (Gu et al., 2024) and REval (Chen et al., 2024) pioneered execution reasoning but are largely confined to isolated, lab-scale code snippets. While tools like CORE (Xie et al., 2025) incorporate static analysis tasks to evaluate reasoning, they lack the structural depth and code breadth (e.g., deep inheritance, cross-module imports) characteristic of large-scale projects. RepoReason bridges this gap by applying deep execution simulation to massive repository scopes, challenging models to maintain logical consistency across extensive file systems.
@@ 2.3 Agentic Software Engineering
At the highest level, SWE-bench (Jimenez et al., 2024) evaluates agents on resolving GitHub issues. However, as an end-to-end evaluation, it acts as a black-box indicator of success or failure. RepoReason complements this by targeting the cognitive root causes of failure. By decoupling reasoning from code editing and utilizing fine-grained metrics (ESV, MCL, DFI), it provides the white-box diagnostic insights required to pinpoint an agent’s specific bottlenecks in context handling or multi-source integration.
@@ 3 RepoReason: Design and Construction
In this section, we delineate the methodological framework of RepoReason, conceptualizing the construction pipeline as a rigorous five-phase defense system. Our objective is to transform authentic software logic into white-box diagnostic tasks while systematically neutralizing systemic vulnerabilities: spurious correlation, rote memorization, and logical superficiality. As illustrated in Figure 1, the pipeline orchestrates sequential gates: (1) Repository Curation to ensure environmental fidelity; (2) Structural Filtering via runtime traces to guarantee logical depth; (3) Execution-Driven Mutation to eliminate parametric leakage; (4) Task Instantiation governed by a deterministic value protocol for stability; and (5) Diagnostic Evaluation establishing fine-grained cognitive profiles. This integration ensures that each task instance necessitates authentic repository-level reasoning and robust state tracking, effectively precluding resolution via shallow pattern matching.
@@ 3.1 Design Philosophy: Assertions as Semantic Anchors
Spurious Correlation in Raw Traces: A straightforward approach might involve randomly querying variable values at specific locations (e.g., “What is the value of variable $x$ at line 10?”). However, this paradigm of raw trace querying is fundamentally flawed for evaluating deep reasoning. It suffers from shortcut learning, as models can often retrieve answers through shallow static analysis of nearby assignment statements, bypassing the need for genuine execution simulation.
@@ 3.2 Foundation: Real-world Repository Curation
To ensure that our semantic anchors are grounded in reality, we curated a set of mature, actively maintained Python repositories. This stage ensures the fidelity of our benchmark, providing the complex, cross-file dependency structures that synthetic datasets lack.
@@ 3.6 Cognitive Diagnostic Framework
To quantify the intrinsic reasoning complexity of a task, we perform dynamic program slicing (Agrawal and Horgan, 1990) rooted at the assertion to obtain the minimal causal computational subgraph along the actual execution path. Dynamic cross-function backward slicing is obtained by tracing back from the assertion along data and control dependencies within a recorded execution trace to isolate the minimal set of statements that causally influence the observed result. Based on this subgraph, we characterize reasoning intensity using three mutually orthogonal metrics rooted in cognitive load theory (Sweller, 1988).
@@ 4.2 Overall Performance: Accuracy across Repositories and Difficulty Levels
1. Explicit Algorithmic Patterns (Cachetools and Toolz): In repositories such as cachetools and toolz, models demonstrate robust and consistent performance, with mean accuracies ranging from 78.00% to 82.00%. These libraries primarily rely on standard algorithmic paradigms—such as cache eviction and functional data flows—characterized by high logical locality and linear execution transitions. The results indicate that LLMs have effectively mastered these procedural patterns, which currently define the cognitive comfort zone for agentic reasoning.
2. Implicit Metaprogramming Logic (Attrs and Jinja2): attrs and jinja2 exhibit a significant performance stratification. Taking attrs as an example, which relies heavily on decorators for implicit class generation (e.g., the dynamic injection of __init__ methods), the performance gap between Claude (72.41%) and Kimi-K2 (43.10%) reveals a fundamental divergence in the internalization of the Python dynamic object model. Frontier models demonstrate the capacity to model runtime behaviors that remain invisible as explicit code lines in static text, whereas other models show a marked deficiency when logic is generated programmatically rather than defined literally.
3. High-Entropy Symbolic and Structural Reasoning (SymPy and NetworkX): sympy and networkx represent the primary reasoning ceiling for current agents. Symbolic transformations in sympy involve abstract, non-linear logical transitions that frequently cause reasoning chains to deviate. At the same time, networkx requires high structured working memory to track complex graph topologies. Claude’s superior performance in networkx (79.77%) confirms its advanced ability to maintain accurate state snapshots during complex traversals without experiencing memory decay or structural confusion.
@@ C.1 Agentic Search Prompting Strategy
C.1.1 Concrete Example: Literal Integer Answer
We present a real prompt instance from the cachetools repository to illustrate the complete Agentic Search prompt structure. This example demonstrates a literal integer answer, where the Agent must infer the value 2 by reasoning about cache operations.
Real Agentic Search Prompt Example Task: Fill in the blank in the Python test code. Blank Assertion: assert ___ == len(cache) Hint: CRITICAL REQUIREMENT: The answer MUST be a literal integer value. IMPORTANT: Return ONLY the literal value, NOT an expression or variable name. Type-specific examples: Boolean: True or False (NOT ’x == y’ or ’bool(x)’) String: "hello" (NOT ’str_var’ or ’x.name’) Number: 42 (NOT ’num_var’ or ’len(x)’) List: [1, 2, 3] (NOT ’list_var’ or ’list(x)’) Dict: {"a": 1} (NOT ’dict_var’ or ’dict(x)’) Instructions: The ’___’ in the assertion represents a missing expression (variable name, function call, or literal). You are a ReadOnlyAgent. You can read files and search the code in the repository to infer what this expression should be. Note: Test function code is not visible in the repository, but you can explore the source code functions called by the test and reason to complete the fill-in-the-blank task. Context: - Test function source code: class LRUCacheTest(unittest.TestCase, CacheTestMixin): Cache = LRUCache def test_update(self): cache = self.Cache(maxsize=2) cache.update({1: 1, 2: 2}) assert ___ == len(cache) # Question assert ___ == cache[1] assert ___ == cache[2] cache.update({1: 1, 2: 2}) assert ___ == len(cache) assert ___ == cache[1] assert ___ == cache[2] cache.update({1: "a", 2: "b"}) assert ___ == len(cache) assert ___ == cache[1] assert ___ == cache[2] - Input parameters: {"self": "<LRUCacheTest>"} Note: The missing expression is likely related to the input parameters or variables defined in the test or the code being tested. Goal: Find the exact expression that should replace ’___’. When you are confident, use the finish tool with the answer. IMPORTANT: The finish tool’s content MUST contain ONLY the answer expression itself, nothing else. Examples: If the answer is a variable: finish with content "result" If the answer is a function call: finish with content "len(items)" If the answer is a string: finish with content ’"hello"’ If the answer is an attribute access: finish with content "C.__attrs__attrs__[1].validator" Do NOT include explanations, reasoning, or any other text in the finish tool content. Only the answer expression.
Analysis. In this example, the Agent must:
  1. 1.
    Understand that cache.update({1: 1, 2: 2}) adds two key-value pairs to a cache with maxsize=2
  2. 2.
    Reason that after the update, len(cache) equals 2
  3. 3.
    Infer that the blank should be filled with the literal integer 2
  4. 4.
    Use the finish tool with content "2" (without quotes in the tool content, as it’s a numeric literal)
This example demonstrates how the prompt guides the Agent to reason about runtime behavior while enforcing strict output constraints for evaluation.